OpenStudy (anonymous):

(Differential equations) I need to solve a PDE (Heat equation) where the initial condition is a constant, T_i. Can I treat the initial condition as a function f(x) in order to solve via separation of variables; and then replace it with T_i after i'm done solving?

OpenStudy (amistre64):

might need to remember the diffyQ for heat :/

OpenStudy (anonymous):

Sure! \[\frac{\partial u}{\partial t}=\alpha \frac{\partial^2 u}{\partial x^2}\]

OpenStudy (amistre64):

soo, to neaten it up a u'' - y' = 0

OpenStudy (amistre64):

... typoed it lol a u'' - u' = 0

OpenStudy (amistre64):

to do separation, you might want to redefine z = u' z' = u'' a dz/du- z = 0 a dz/du = z dz/z = du/a ln(z) = u/a + C z = Ce^(u/a) = u'

OpenStudy (anonymous):

Don't forget that u is a function of two variables, not one

OpenStudy (amistre64):

.... doh!! those are partials ;/

OpenStudy (anonymous):

hehe

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/DE/TheHeatEquation.aspx then thisll be more congent than i can be :)

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/DE/SeparationofVariables.aspx and this is the "later one" for the separation of variables solution to it

OpenStudy (anonymous):

Looks like yes, one of his examples did exactly what I thought; treating the constant initial condition as a function of x, allowing separation of variables. thanks.

OpenStudy (amistre64):

:) good luck